The resistance R, is a variable resistor that can take on values in the range 0 R,24 Find the value of R1 = R1a that maximize Find i1a , R1a , v1b , R1b , P1c , and R1c
Answer #1

1 26 R. = Ria=on them of maximon currow will flowing ON 7. 24 (on means shar circuit) 12- 24 Ics24.26 = By KCL I, a= 2 + 4x +

31 Open R, q find Voc = 4th Shoa Va vo 122 240 @ zur va=240 Apply nodal analynn at vi 18 +55 +2 = 0 25+ b +7L=0 Vo = -240 in

If you have any doubt please make a comment. Thank you